OpenStudy (anonymous):

A horse races once around a circular track in a time of 106 s, with a speed of 31 m/s. What is the radius of the track? A horse races once around a circular track in a time of 106 s, with a speed of 31 m/s. What is the radius of the track? @Physics

OpenStudy (anonymous):

\[2\pi*r=31*106\] r = 523 m

OpenStudy (anonymous):

T = 106 s T = 2 * pi / w w = 2* pi / T = 2 * 3.14 / 106 = 0.05927 rad/s and v = r * w ; r = v / w r = 31/ 0.05927 r = 523 m